OpenStudy (marigirl):

check my ans please: I am working on a differentiation question. Can you please check if i have differentiated correctly below and solved for x

OpenStudy (marigirl):

\[A=4x-2x^2-\frac{ \pi x^2 }{ 2 }\] \[A'=4-4x-\pi x\]

OpenStudy (michele_laino):

correct!

OpenStudy (marigirl):

now to set f'(x)=0 \[4-4x-\pi x=0\] \[x=\frac{ 4 }{ 4+ \pi }\]

OpenStudy (marigirl):

so then \[2x=2*\frac{ 4 }{ 4+ \pi }\] \[=\frac{ 8 }{ 4+\pi }\]

OpenStudy (michele_laino):

correct!

OpenStudy (marigirl):

then i have \[h=2-x-\frac{ \pi x }{ 2 }\]

OpenStudy (marigirl):

no sorry, that is my h . i need to sub in x

OpenStudy (marigirl):

then i have: \[h=2-\frac{ 4 }{ 4+\pi }-\frac{ \pi }{ 2 }\frac{ 4 }{ 4+x }\]

OpenStudy (marigirl):

\[h=2-\frac{ 4 }{ 4+\pi }-\frac{ 4 \pi }{ 8+2 \pi }\]

OpenStudy (marigirl):

\[h=2-\frac{ 4 }{ 4+\pi }-\frac{ 2\pi }{ 4+\pi }\]

OpenStudy (michele_laino):

is \(A(x)\) the area of a triangle?

OpenStudy (marigirl):

\[h=2-\frac{ 4-2\pi }{ 4+\pi }\]

OpenStudy (marigirl):

my area equation and h equation is both correct

OpenStudy (marigirl):

\[h=\frac{ 2(4+\pi) }{ (4+\pi) }-\frac{ 4-2\pi }{ 4+\pi }\] \[h=\frac{ 8+2\pi }{ 4+\pi }-\frac{ 4-2\pi }{ 4+\pi }\]

OpenStudy (marigirl):

\[h=\frac{ 4 }{ 4+\pi }\]

OpenStudy (marigirl):

the answer reads (8/4+pi) and i got 4 .. .

OpenStudy (michele_laino):

from your drawing, I can write the area \(A(x)\) as below: \[A\left( x \right) = 2hx + \frac{{\pi {x^2}}}{2}\]

OpenStudy (marigirl):

this is the actual question

OpenStudy (michele_laino):

\(h(x)\) is correct!

OpenStudy (marigirl):

maybe it is an error in the model answers?

OpenStudy (michele_laino):

also \(A(x)\) is correct!

OpenStudy (michele_laino):

maximum of \(A(x)\) is at: \[x = \frac{4}{{4 + \pi }}\] so, correct!

OpenStudy (marigirl):

great thanks so much. i did about 4 pages of work trying to see where i went wrong..

OpenStudy (michele_laino):

by substitution into \(h(x)\), I get: \[{x_0} = \frac{4}{{4 + \pi }},\quad h\left( {{x_0}} \right) = \frac{8}{{2\left( {4 + \pi } \right)}}\]

OpenStudy (marigirl):

where am i going wrong

OpenStudy (michele_laino):

namely: \[h\left( {{x_0}} \right) = \frac{8}{{2\left( {4 + \pi } \right)}} = \frac{4}{{4 + \pi }}\]

OpenStudy (michele_laino):

of, course the base is: \[2{x_0} = \frac{8}{{4 + \pi }}\]

OpenStudy (marigirl):

so maybe a printing error in the book?

OpenStudy (michele_laino):

I think it is possible

OpenStudy (marigirl):

could you look over another question plz? its the wording of it i am not really sure of

OpenStudy (michele_laino):

ok!

OpenStudy (marigirl):

last question

OpenStudy (michele_laino):

please wait, I'm working on it...

OpenStudy (marigirl):

sure thanks..

OpenStudy (marigirl):

could u also explain what part (c) is actually asking me to do please?

OpenStudy (michele_laino):

I start from point c)  the area of the rectangle is: \[A\left( x \right) = 2x \cdot \left( {9 - {x^2}} \right)\]